### v3

@@ Abstract
Deep learning has yielded remarkable outcomes in various domains. However, the challenge of requiring large-scale labeled samples still persists in deep learning. Thus, data augmentation has been introduced as a critical strategy to train deep learning models. However, data augmentation suffers from information loss and poor performance in small sample environments. To overcome these drawbacks, we propose a feature augmentation method based on shape space theory, i.e., feature augmentation on Geodesic curve, called FAGC in brevity. First, we extract features from the image with the neural network model. Then, the multiple image features are projected into a pre-shape space as features. In the pre-shape space, a Geodesic curve is built to fit the features. Finally, the many generated features on the Geodesic curve are used to train the various machine learning models. The FAGC module can be seamlessly integrated with most machine learning methods. And the proposed method is simple, effective and insensitive for the small sample datasets. Several examples demonstrate that the FAGC method can greatly improve the performance of the data preprocessing model in a small sample environment.
keywords:
[…]
@@ 1 Introduction
As a powerful artificial intelligence technology, deep learning has made revolutionary breakthroughs in computer vision, natural language processing, speech recognition, and other related fields [1, 2, 3]. The training of deep learning models usually requires large amounts of labeled data to achieve excellent performance, but it is difficult to obtain large-scale and high-quality labeled data. To fully exploit the information in the small sample, data augmentation is introduced as an essential strategy to compensate for the lack of limited labeled data in training deep learning models. Data augmentation provides more learning opportunities to improve the performance and generalization of the deep learning models by expanding the training dataset and increasing the diversity of the data.
There are three main types of data augmentation methods. The first one is to use perturbation images, which add some factors to the original image, such as panning, rotating, flipping, cropping, adding noise, and other image processing methods [4]. These simple methods have some disadvantages. First, the distribution of the perturbation images ought to be close to the distribution of the original images. Furthermore, some perturbation image methods, such as translation and rotation, suffer from the padding effect. After the operation is completed, part of the image is moved out of the boundary and lost. Thus, some interpolation methods can select one of the four pixels around the blank position and fill it with the blank section [5]. These processing methods can weaken the model’s attention to the critical regions and affect the model’s performance. The current study tends more toward sophisticated methods, such as Random Erasure, Mixup [6], Cutout [7], RandAugment [8], etc. These methods can improve the learning ability and generalization ability of the models of deep learning. Another type of data augmentation is generative models, which generate the virtual data from the real data, such as GANs [9] and VAEs [10]. These models have been widely used to their excellent generative capabilities. However, the generative models still rely on large data sizes, leading to the generally poor quality of images generated by these generative models in the face of limited datasets.
The third type of data augmentation method is to augment the features extracted from the data, i.e., feature augmentation. The key idea is to reshape the decision boundaries of rare categories by directly manipulating the data in the feature space and enhancing the representational ability of the data. For example, DeVries and Taylor proposed a data augmentation approach based on feature space [11]. Because they found that when traversing along a manifold, the probability of encountering real samples in the feature space is higher than in the input space. The SMOTE method selects a synthetic sample point by randomly choosing a point on the line that intersects k similar samples in feature space [12]. The DBSMOTE algorithm [13], as a derivative of SMOTE, generates new synthetic samples by calculating the distances from the cluster centers. Although the generated samples are very close to the original sample distributions, it also increases the risk of overfitting. Recently, Vikas Verma et al. proposed a new method, called manifold mixup, for training neural networks on interpolation of hidden representations [14]. The method encourages the neural network to perform uncertainty analysis of the feature space, which was not visible during the training phase and, in turn, generates more discriminative features. ISDA enhances the data samples by transforming the CNN features along the semantic directions [15]. However, these methods always result in decreased performance and increased complexity when utilized with limited datasets.
In order to solve the shortcomings of the feature augmentation methods, this paper proposes FAGC, a feature augmentation method on Geodesic curve based on the shape space theory. The shape space theory [16, 17, 18], as a kind of manifold theory, mainly focuses on the geometric attributes, structure, and arrangement of the data. The theory has been widely used in image processing due to its remarkable interpretability of data and insensitivity to limited datasets [19, 20, 21, 22]. For example, Kilian et al. proposed to use shape space for sampling and projection on a highly sparse collection of three-dimensional models to generate geometrically and semantically valid samples [23]. Han et al., on the other hand, proposed an object recognition method using the shape space theory. They projected object contours into shape space and constructed Geodesic curves consistent with the diverse potential shapes of a given object type [24]. In this paper, we extract features from images by a deep learning model and then project these image features into a pre-shape space. Then, the Geodesic curve fits these features and generates new features for machine learning or deep learning training. Thus, FAGC achieves data augmentation in a small sample scenario.
@@
A concise overview of our framework is presented in Figure 1, highlighting our primary contributions as follows:
  • The FAGC module can generate the data flexibly and conveniently in the small sample environments.
  • A Geodesic curve is constructed to fit the projected images into pre-shape space.
  • The random probability and the influence factors are introduced in the classifier module for downstream tasks.
  • Empirical results demonstrate that FAGC significantly improves the performance of downstream tasks on different benchmark datasets.
Figure 1: An overview of FAGC. The image data sequentially passes through the feature extraction module, the expand module, and the project module, which converts the image data into features and projects them into the pre-shape space. In the pre-shape space, the Geodesic curves are constructed by minimizing the sum of squares of Geodesic distance. The Geodesic curves corresponding to each category are constructed. Then, new features are generated according to the categories and given to the machine learning models to achieve the effect of feature augmentation.
@@ 2.1 Pixel-level augmentation
Data augmentation has an important role in the evolution of machine learning. The earliest demonstrations of the effectiveness of data augmentation came from simple transformations such as horizontal flipping of pixels, color augmentation, and random cropping. Most of these techniques operate directly on the image and are easy to implement. Then, image augmentation methods based on image erasure, in which the main idea is to remove one or more sub-regions of an image and replace the pixel values of these sub-regions with constant or random values, were proposed. DeVries et al. proposed a simple regularisation technique called Cutout, whereby a square region of the input is randomly masked during the training of a convolutional neural network [7]. The technique improves the robustness and overall performance of the convolutional network. Singh et al. proposed Hide-and-Seek to randomly hide blocks in the training image, which forces the network to look for other relevant content while the most discriminative content is hidden [25]. Zhong et al. proposed random erasure, where a randomly selected rectangular region in the image replaces other pixels with random values [26]. In addition, some image-mixing methods mainly achieve augmentation by mixing two or more images or sub-regions of an image into a new image. Zhang et al. proposed a general synthesis method, Mixup [6]. Instead of averaging the intensities of two images, Mixup performs a convex combination of sample pairs and their labels. As a result, Mixup establishes a linear relationship between data augmentation and supervised labels and can regularize neural networks. Drawing on the idea of Mixup, Yun et al. proposed CutMix [27], which replaces the removed region with another image or images instead of simply removing pixels from the training set or mixing the images to produce a more natural image than Mixup. One obvious drawback of the augmentation technique, however, is that the mixed image itself does not make any sense from a human perspective. The performance gains from mixed images are also difficult to understand and explain. One possible explanation is that the size of the dataset enhances the ability to characterize low-level features (e.g., lines and edges).
@@ 2.2 Feature-level augmentation
Feature augmentation is mainly a method of transforming and enhancing data in feature space. In the face of image data, the traditional methods mainly focus on image extraction by feature operators to obtain more valuable feature representations to feed machine learning models. For example, SIFT [28], HOG [29], and LBP [30] feature operators can extract useful feature information from images. However, it has many drawbacks, such as the need to manually design the feature extractor and poor performance in complex scenes. With the development of deep learning, convolutional neural networks can automatically learn image feature representations without manually designing feature extractors and show more powerful generalization ability and performance. Therefore, more and more image processing methods are inclined to use convolutional neural networks for feature extraction. As an example, the stochastic feature-enhanced SFA algorithm utilizes class-conditional covariance matrix information during neural network training to improve model performance [31]. In addition, Li et al. introduced a method of implicit feature enhancement known as moment exchange [32]. The method encourages the model to use the moment information of potential features at another moment to replace the current moment information, in order to achieve the effect of feature enhancement. VAE is also suitable for feature augmentation because the decoder can reconstruct the low-dimensional representations extracted by the encoder into the original image [10]. Similarly, it is possible to augment the feature representation in a convolutional neural network in feature space by changing the output layer of the network so that the output is a low-dimensional feature instead of a category label. These features representations can train machine learning models such as plain Bayes, support vector machines, and fully connected layer networks [33], thereby improving model performance for downstream tasks. In our proposed feature augmentation method, the feature extraction process also draws on the idea, using convolutional neural networks for feature extraction, and then a series of augmentation processes are performed on the extracted features.
@@ 2.3 The shape space theory
The theory of shape space, firstly introduced by Kendall, is used to describe an object and all its equivalent variations in a non-Euclidean space [16]. In shape space, the position, scale, and orientation of a shape can be neglected. The shape space theory is often used in object recognition scenarios, where the distance between objects in shape space determines the outcome of object recognition. The feature representing the shapes of the objects are first projected into the pre-shape space, i.e., all the shapes of the features of the $p$ coordinate points are embedded in a unit hyper-sphere [17, 34, 35], denoted as $S_{*}^{2p-3}$. Any shape is a point or vector on this hyper-sphere, and all changes in the shape, i.e., position, scale scaling, and 2D rotation, result in a new shape that lies on a great circle in $S_{*}^{2p-3}$, denoted $O(v)$, where $v$ denotes that the shape is on $S_{*}^{2p-3}$. The set of all shapes in the unit hyper-sphere $S_{*}^{2p-3}$ forms the orbit space $\sum_{2}^{p}$, which is a shape space described as follows:
$\sum{\begin{subarray}{c}p\\[5.0pt] 2\end{subarray}}=\{O(v):v\in S_{*}^{2p-3}\},$ (1)
An example of pre-shape space and shape space is shown in Figure  2, where pre-shapes of the same polygon at different locations, scales, and orientations are all on the same great circle of $S_{*}^{2p-3}$, where the great circle serves as a point or vector in the shape space. Rather than the Euclidean distance, the Geodesic distance is the shortest distance between two points in the pre-shape space $S_{*}^{2p-3}$. Assuming that $v_{1}$ and $v_{2}$ are two features in the pre-shape space, the Geodesic distance between $v_{1}$ and $v_{2}$ is defined as [36]:
$d(v_{1},v_{2})=cos^{-1}(\langle v_{1},v_{2}\rangle),$ (2)
@@
$d_{p}[O(v_{1}),O(v_{2})]=\text{inf}[\cos^{-1}(\langle\alpha,\beta\rangle):\alpha\in O(v_{1}),\beta\in O(v_{2})],$ (3)
However, the $d_{p}$ calculation in the formula is very complicated and is usually projected into the complex domain space so that Formula (3) can be described by the Procrustean distance [39]:
@@
In pre-shape space, points on a Geodesic curve can be used to describe a set of changing shapes. Evans et al. proposed a method to fit the data in the pre-shape space by obtaining a set of orthogonal points in the pre-shape space to construct the Geodesic curve [42, 43]. Based on this, Han et al. proposed a method to construct Geodesic curves and thus expand the data in the pre-shape space using some small amount of data, where the two endpoints of the curves can be non-orthogonal [24]. Assuming that $v_{1}$ and $v_{2}$ denote two points in the pre-shape space, the Geodesic curve formula can be derived as follows:
@@
where $\theta_{(v_{1},v_{2})}$ denotes the Geodesic distance between $v_{1}$ and $v_{2}$, and $s$ is the angle that denotes the Geodesic distance between the new point $v$ and $v_{1}$. If $0\leq s\leq\theta_{(v_{1},v_{2})}$, then $v_{1}\leq\Gamma_{(v_{1},v_{2})}\leq v_{2}$. Following the path, a series of new data or points can be generated in the pre-shape space, and all of these points are equivalent variations between $v_{1}$ and $v_{2}$, which is extremely helpful for expanding the size of the dataset. Based on Formula (5), the work constructs a Geodesic curve in the pre-shape space to fit the features of multiple images and generates new features along the curve to achieve feature augmentation.
@@ 3 Methods
The FAGC method consists of four steps: image feature extraction, projection of the image features into the pre-shape space as features, construction of a Geodesic curve to fit the features and generation of new data along the Geodesic curve. The main steps outlined above are depicted in Figure 3.
@@ 3.1 Image feature extraction
The initial step of the FAGC method involves mapping the input data from the pixel space to the pre-shape space. Typically, the quality of features is significantly influenced by the feature extraction method. Here, we adopt the deep learning method. Complex neural network models have a tendency to extract highly discriminative features, and the utilization of high-quality features can significantly improve the performance of downstream tasks. The image features can be extracted from each layer of the neural network models. The higher layers of the network often exhibit a stronger capability to learn more intricate and comprehensive feature representations. In this paper, we employ a Vision Transformer (ViT) network [44] of which the final fully connected layer is removed as the feature extractor. The samples are fed into the pre-trained ViT network for fine-tuning, and the fine-tuning process is continued until the cross-entropy loss stabilizes. The output features from the trained ViT network represent the features in the feature space.

### v4

@@ Abstract
Due to the constraints on model performance imposed by the size of the training data, data augmentation has become an essential technique in deep learning. However, most existing data augmentation methods are affected by information loss and perform poorly in small-sample scenarios, which limits their application. To overcome the limitation, we propose a Feature Augmentation method on Geodesic Curve in the pre-shape space, called the FAGC. First, a pre-trained neural network model is employed to extract features from the input images. Then, the image features as a vector is projected into the pre-shape space by removing its position and scale information. In the pre-shape space, an optimal Geodesic curve is constructed to fit the feature vectors. Finally, new feature vectors are generated for model learning by interpolating along the constructed Geodesic curve. We conducted extensive experiments to demonstrate the effectiveness and versatility of the FAGC. The results demonstrate that applying the FAGC to deep learning or machine learning methods can significantly improve their performance in small-sample tasks.
Keywords: Feature augmentation, Shape space, Pre-shape space, The Geodesic curve
@@ 1 INTRODUCTION
As a powerful artificial intelligence technology, deep learning has made revolutionary breakthroughs in computer vision, natural language processing, speech recognition, and other fields [1, 2, 3]. A well-performing deep learning model typically requires large-scale, high-quality labeled data for training, but it is often not available in reality. The small-sample problem has become a major factor impeding the broader application of deep learning methods in certain real-world scenarios. For instance, in the field of materials science, acquiring a single SEM image of a copper alloy is often costly, potentially amounting to several hundred dollars. The key to addressing the small-sample problem lies in providing sufficient knowledge representation from limited samples, thereby enhancing its generalization capability.
Some novel approaches addressing the small-sample challenge have recently emerged, showing significant performance gains across multiple benchmarks. Transfer learning is a mainstream approach in deep learning to address the problem of data scarcity [4]. Its core idea is to transfer knowledge obtained from large-scale datasets to the training process of the target dataset, mitigating the lack of information. In practical applications, transfer learning most commonly follows a two-stage training paradigm, i.e., pre-training and fine-tuning. The transfer model first learns generic feature extraction capabilities on a large-scale dataset, and is subsequently fine-tuned on the target dataset to adapt to specific tasks. However, the methodology necessitates a large-scale source dataset relevant to the target domain, which often unmet in practical scenarios. Crucially, transfer performance is highly dependent on the inter-dataset relevance. Leveraging transfer learning strategies, meta-learning-based Few-Shot Learning (FSL) have been proposed to address the small-sample problem [5]. The FSL first acquires a similarity function by training on a large-scale base dataset rich in category and sample information, developing the model’s ability to adapt to novel categories. During inference, the FSL makes final predictions by computing pairwise similarity measures between query set instances and support set prototypes. Although the FSL solutions are generally applicable and can adapt to most few-shot target datasets, they often prove less efficient when the target task is well-defined. By contrast, data augmentation offers a more straightforward solution to small-sample problems through artificial expansion of limited datasets [6].
While conventional augmentation strategies (e.g., geometric transformations or color jittering) offer baseline improvements, they often fail to capture complex data distributions, especially in small-sample scenarios. With the advancement of deep learning, data augmentation methodologies have evolved along two main directions by leveraging the powerful representational capabilities of large-scale models. One approach is to leverage deep generative techniques such as GAN [7], VAE [8], and Diffusion [9] models to synthesize virtual samples that conform to the real data distribution for training. While these virtual data enhance sample diversity, developing effective generative models under small-sample constraints remains a significant challenge. In addition, in some fields with high security requirements, such as medical applications, the generated virtual data lacks sufficient reliability. Another approach is to directly use large models to extract image features, and then data augmentation is achieved at the feature level. In Ref [10], the probability of encountering real samples in the feature space is higher than that in the input space when traversing along the manifold. Despite boosting downstream task performance, current feature augmentation techniques remain perform poorly in small-sample settings.
To compensate for this deficiency, we introduce the shape space theory to propose Feature Augmentation on Geodesic Curve (FAGC) in pre-shape space, delivering effective feature augmentation under data scarcity. The shape space theory [11, 12, 13], as a kind of manifold theory, mainly focuses on the geometric attributes and structures of the data. The theory has been widely used in image processing due to its remarkable interpretability of data and insensitivity to limited datasets [14, 15, 16, 17]. As demonstrated in [18], the pre-shape space exhibits inherent robustness to small-sample data and enables seamless integration with deep learning models. Furthermore, as indicated in [19], the Geodesic curves in the pre-shape space contain rich geometric transformation information that can improve object recognition accuracy. Building upon the current research foundation, the FAGC provides a comprehensive manifold feature augmentation framework for machine learning pipelines. The FAGC first represents image data as feature vectors within the pre-shape space, involving processes such as feature extraction, modeling, and projection. The core of the FAGC lies in its designed algorithm, which constructs a optimal Geodesic curve in the pre-shape space that fits multiple initial feature vectors. The curve minimizes the sum of Geodesic distances to all initial feature vectors under the metric of the pre-shape space. Then, new feature vectors are sampled along the curve to achieve feature augmentation. Finally, to effectively utilize both the augmented data and the real data during downstream task training, a corresponding loss function construction method is formulated.
@@
Figure 1 provides a concise overview of our framework. Our main contributions are summarized as follows:
  • A feature augmentation method called the FAGC is proposed. The FAGC achieves efficient feature augmentation under small-sample conditions by constructing the optimized Geodesic curves in the pre-shape space.
  • The FAGC can be seamlessly integrated into the machine learning methods or deep learning methods, demonstrating its strong performance in extensive small-sample scenarios, particularly when data is extremely scarce.
  • Multiple ablation studies bridge the theoretical foundation of the FAGC to its practical applications. The visual experiments further validate the safety and reliability of the data augmented by the FAGC, expanding its applicability scope.
Figure 1: An overview of the FAGC method.
@@ 2.1 Pixel-level augmentation
The effectiveness of data augmentation was first demonstrated through simple pixel-level transformations, such as horizontal flipping, color enhancement, and random cropping [20]. These pixel-level methods have some disadvantages. First, the distribution of the perturbation images ought to be close to the distribution of the original images. Furthermore, some perturbation image methods, such as translation and rotation, suffer from the padding effect. After the operation is completed, part of the image is moved out of the boundary and lost. Thus, some interpolation methods select one of the four pixels around the blank position and use it to fill the blank area [21]. These processing methods weaken the model’s attention to the critical regions and affect the model’s final performance. Recent studies have placed greater emphasis on sophisticated methods, such as Random Erasure [22], Mixup [23], Cutout [24], RandAugment [25], etc. These methods effectively increase the amount of training data, but the improvement in diversity is limited. With the advancement of large-scale models, data-driven pixel-level image reconstruction has been leveraged for data augmentation. For example, AutoMix [26] proposes an automated mixing framework that enables models to generate mixed samples conducive to classification performance. Another reconstruction method is generative modeling, which generates the realistic image from the real image, such as GAN [7], VAE [8], and Diffusion [9]. These models have been widely used due to their excellent generative capabilities. However, the quality of generative models still heavily depends on the amount of data, often resulting in poor image quality when the training data is limited.
@@ 2.2 Feature-level augmentation
Feature-level augmentation primarily refers to the method of transforming and augmenting data in the feature space. Extracting distinctive features from images is a prerequisite for feature augmentation. Early methods typically relied on manually designed feature operators, such as SIFT [27], HOG [28], and LBP [29], which were often less effective. With the advancement of deep learning, deep models have been employed to replace these manual approaches for extracting high-quality features. Consequently, feature augmentation has gained significant research attention relying on deep learning techniques. For example, the stochastic feature-augmented SFA algorithm expands data by randomly applying noise information to the output of the intermediate layer during model training [30]. In [31], an implicit feature augmentation method called moment exchange is proposed. The method encourages the model to replace the current moment information with the moment information of latent features at another time. There are also some feature augmentation methods that are improvements upon existing input-space data augmentation techniques. As in [32], SMOTE is applied at the feature level, enhancing the accuracy of imbalance fault diagnosis in gas turbines. Similarly in [33], feature-level Mixup techniques, when applied to intermediate neural network layers, have successfully enhanced model performance. Although existing feature augmentation methods have demonstrated effectiveness, further improving model performance in small-sample scenarios remains a significant challenge. In this work, we leverage the shape space theory, which has been proven effective for small-sample scenarios [18], to enable efficient feature augmentation in the manifold space.
@@ 2.3 The shape space theory
The shape space theory, firstly introduced by Kendall, is used to describe an object and all its equivalent variations in a non-Euclidean space [11]. In the shape space, the position, scale, and orientation of a shape can be neglected. The shape space theory is often used in object recognition scenarios, where the distance between objects in shape space determines the outcome of object recognition. The feature representing the shapes of the objects are first projected into the pre-shape space, i.e., all the shapes of the features of the $p$ coordinate points are embedded in a unit hyper-sphere [12, 34], denoted as $S_{*}^{2p-3}$. Any shape is a point or vector on this hyper-sphere, and all changes in the shape, i.e., position, scale scaling, and 2D rotation, result in a new shape that lies on a great circle in $S_{*}^{2p-3}$, denoted $O(v)$, where $v$ denotes that the shape is on $S_{*}^{2p-3}$. The set of all shapes in the unit hyper-sphere $S_{*}^{2p-3}$ forms the orbit space $\sum_{2}^{p}$, which is a shape space described as follows:
$\sum{\begin{subarray}{c}p\\[5.0pt] 2\end{subarray}}=\{O(v):v\in S_{*}^{2p-3}\}.$ (1)
An example of pre-shape space and shape space is shown in Figure 2. Pre-shapes of the same polygon at different locations, scales, and orientations are all on the same great circle of $S_{*}^{2p-3}$, where the great circle serves as a point or vector in the shape space. Rather than the Euclidean distance, the Geodesic distance is the shortest distance between two points in the pre-shape space $S_{*}^{2p-3}$. Assuming that $v_{1}$ and $v_{2}$ are two features in the pre-shape space, the Geodesic distance between $v_{1}$ and $v_{2}$ is defined as [19]:
$d(v_{1},v_{2})=\cos^{-1}(\langle v_{1},v_{2}\rangle),$ (2)
@@
$d_{p}[O(v_{1}),O(v_{2})]=\text{inf}[\cos^{-1}(\langle\alpha,\beta\rangle):\alpha\in O(v_{1}),\beta\in O(v_{2})].$ (3)
However, the $d_{p}$ calculation in the formula is very complicated and is usually projected into the complex domain space. Therefore, Formula (3) can be described by the Procrustean distance [36]:
@@
In the pre-shape space, points on a Geodesic curve can be used to describe a set of changing shapes. Evans et al. proposed a method to fit the data in the pre-shape space by obtaining a set of orthogonal points in the pre-shape space to construct the Geodesic curve [39, 40]. Based on this, Han et al. proposed a method to construct the Geodesic curve and thus expand the data in the pre-shape space using some small amount of data, where the two endpoints of the curves can be non-orthogonal [41]. Assuming that $v_{1}$ and $v_{2}$ denote two points in the pre-shape space, the Geodesic curve formula can be derived as follows:
@@
where $\theta_{(v_{1},v_{2})}$ denotes the Geodesic distance between $v_{1}$ and $v_{2}$, and $s$ is the angle that denotes the Geodesic distance between the new point $v$ and $v_{1}$. If $0\leq s\leq\theta_{(v_{1},v_{2})}$, then $v_{1}\leq\Gamma_{(v_{1},v_{2})}\leq v_{2}$. Following the curve, a series of new data or points can be generated in the pre-shape space, and all of these points are equivalent variations between $v_{1}$ and $v_{2}$, which is extremely helpful for expanding the size of the dataset.
The shape space theory has been applied to various vision-related tasks, such as action recognition [42], object recognition [41], and point cloud registration [43]. Existing methods typically treat two pre-shapes as fixed endpoints to construct Geodesic curves for data augmentation. These generated samples encapsulate rich latent transformation information between two pre-shapes. In practical scenarios, however, multiple pre-shapes often coexist, rendering existing pairwise Geodesic construction methods inadequate for leveraging holistic dataset information. To address this limitation, we propose a more practical method for Geodesic curve construction among multiple pre-shapes.
@@ 3.1 Image feature extraction
The initial step of the FAGC involves mapping the input data from the pixel space to the pre-shape space. Typically, the quality of features is significantly influenced by the feature extraction method. Here, we adopt the deep learning method. Complex neural network models have a tendency to extract highly discriminative features, and the utilization of high-quality features can significantly improve the performance of downstream tasks. The image features can be extracted from each layer of the neural network models. The higher layers of the network often exhibit a stronger capability to learn more intricate and comprehensive feature representations. In this paper, we employ a Vision Transformer (ViT) network [44] of which the final fully connected layer is removed as the feature extractor. The samples are fed into the pre-trained ViT network for fine-tuning, and the fine-tuning process is continued until the cross-entropy loss stabilizes. The output features from the trained ViT network represent the features in the feature space.